You need to produce a buffer solution that has a pH of 5.50. You already have a solution that contains 10. mmol (millimoles) of acetic acid. How many millimoles of acetate (the conjugate base of acetic acid) will you need to add to this solution? The pKa of acetic acid is 4.74.
Answer #1

For a buffer, we know that:

apply Henderson Hasselbach Equation

pH = pKa + log(conjugate base / weak acid)

substitute known data

5.5 = 4.74 + log(conjugate base / weak acid)

conjugate base / weak acid = 10^(5.5-4.74) = 5.754399

now..

relate:

HA <-> H+ + A-

mmol of HA inintially = 10

mmol of acetate required = ?

conjugate base / 10 = 10^(5.5-4.74) = 5.754399

conjugate base = 10*5.754399 = 57.54 mmol of actate required